# Perceptions to Beliefs: Exploring Precursory Inferences for Theory of Mind in Large Language Models

### v2

@@ 6 Related Work
Methods for Improving LLM’s Theory of Mind
Previous research has explored several methods to enhance LLM’s ToM ability. SymbolicToM (Sclar et al., 2023) tracks multiple characters’ beliefs using graphical representation to provide LLMs the context in the target character’s point of view. However, the necessity to construct the belief state graph restricts its adaptability in complex scenarios involving diverse relationships and interactions between entities. ToM-LM (Tang and Belle, 2024) improves performance through LLM fine-tuning, while it requires additional training resources. SimToM (Wilf et al., 2023) improves LLM’s ToM ability through prompt tuning and highlights the significance of perspective-taking.
@@
Table 2: Performance comparison of perception-to-belief inference and PercepToM+Oracle.
@@ 7 Conclusion
Inspired by the psychology literature, we evaluated the precursory inferences for human theory of mind (ToM) in large language models (LLM) aiming to broaden our insight into their ToM capabilities. To this end, we constructed Percept-ToMi and Percept-FANToM, perception-augmented ToM benchmarks by annotating character perceptions about the contexts. Through evaluations and analyses on eight state-of-the-art LLMs, we found that they perform reasonably well in inferring others’ perceptions but struggle with inferring others’ beliefs based on that perceptual information. Based on these findings, we proposed a new framework, PercepToM, for improving LLM’s ToM reasoning. Our framework leverages LLMs’ strength in perception inference and enhances their perception-to-belief inference by extracting the relevant contexts. We expect our work to provide insights and enable further in-depth studies into the extent of LLMs’ ToM capabilities and targeted improvements in their weaknesses.
@@ 8 Limitations
We compare LLMs’ ToM performances between true belief and false belief scenarios, but not those between the different orders of ToM questions (e.g., first-order and second-order). Since higher-order ToM requires more inference steps, it will be also interesting to examine the differences in model behavior and capability in solving different orders of ToM questions in future work.
@@ 9 Societal and Ethical Considerations
Our use of FANToM dataset is consistent with its intended use, which is evaluation. We have adhered to the licenses of the benchmarks, ToMi and FANToM, in processing them to create our benchmarks, Percept-ToMi and Percept-FANToM. We plan to make our benchmarks publicly available with the license of Attribution-Noncommercial 4.0 International (CC BY-NC 4.0), allowing sharing and adapting of the material.
@@ A.1 Manual Verification of Perception Annotation in Percept-ToMi
Through manual verification of the perceiver annotations in Percept-ToMi generated by SymbolicToM, we modify some of them. First of all, the perceiver of distractor sentences in ToMi, which describe a character’s opinion about an object, should be the character holding the opinion. However, the SymbolicToM-generated perceiver annotation also includes other characters. We therefore correct the perceivers for all distractor sentences.
The sentences preceding the location-disambiguating sentence, which specifies object locations, were annotated with ‘none’ as the perceiver. We align the perceiver annotations of these sentences with those of the subsequent location-disambiguating sentence, since they are always paired and have the same perceivers. The example perceiver annotations corrected by manual verification are shown in Table 3.
@@
A.2 Perception Annotation Criteria of Percept-FANToM
The criteria for annotating perceivers in the FANToM dataset are as follows.
  1. 1.
@@ A.2 Perception Annotation Criteria of Percept-FANToM
Appendix B Prompt Examples
@@ B.1 Perception Inference
B.2 Perception-to-Belief Inference
@@
Appendix C Input and Output Examples of PercepToM Pipeline
@@
[Input Prompt]: Story: Ella likes the suit. Ella entered the cellar. Lucas entered the cellar. Benjamin entered the porch. The boots is in the cupboard. The cupboard is in the cellar. Lucas exited the cellar. Benjamin exited the porch. Ella likes the sweatshirt. Lucas entered the porch. Ella moved the boots to the pantry. The pantry is in the cellar.
@@
[Input Prompt]: Gianna: Guys, I’ve really enjoyed sharing our pet stories, but I need to excuse myself. I need to change clothes for a meeting later. Talk to you later!
@@
[Out]: ["Ella likes the suit.": ["Ella"],
@@
[Out]: [{"Gianna: Guys, I’ve really enjoyed sharing our pet stories, but I need to excuse myself. I need to change clothes for a meeting later. Talk to you later!": ["Gianna", "Sara", "Javier"]},
@@
[Out]: Lucas entered the cellar. The boots is in the cupboard. The cupboard is in the cellar. Lucas exited the cellar. Lucas entered the porch.
@@
[Out]: Gianna: Guys, I’ve really enjoyed sharing our pet stories, but I need to excuse myself. I need to change clothes for a meeting later. Talk to you later!
@@
[Input Prompt]: Here are the past scenes in sequence that Lucas knows about.
@@
[Input Prompt]: Here are the past utterances in sequence that Gianna is aware of.
@@
Appendix D LLM Performances on Percept-ToMi and Percept-FANToM
@@
Appendix E Performance Comparison Between PercepToM and SymbolicToM

### v3

@@ 5 Related Work
Table 2: LLMs perform significantly better in PercepToM+Oracle than perception-to-belief inference across most scenarios. Since the only difference between the two tasks is the inclusion of a perspective context extraction step in PercepToM+Oracle, the result suggests that LLMs struggle to suppress irrelevant perception information when solving ToM questions.
Improving LLM’s Theory of Mind
Previous research has explored several methods to enhance LLM’s ToM ability. SymbolicToM (Sclar et al., 2023) tracks multiple characters’ beliefs using graphical representation to provide LLMs the context in the target character’s point of view. However, the necessity to construct the belief state graph restricts its adaptability in complex scenarios involving diverse relationships and interactions between entities. SimToM (Wilf et al., 2023) improves LLM’s ToM ability through prompt tuning and highlights the significance of perspective-taking. ToM-LM (Tang and Belle, 2024) improves performance through LLM fine-tuning, while it requires additional training resources.
While SymbolicToM and SimToM achieve high performance on ToMi, their algorithms are tailored for the ToMi structure. For example, SimToM’s prompts include ToMi-specific hints, such as specific lists of events a character should be aware of for ToMi and instructions for output to depend on a fixed pattern in BigToM stories. As a result, SimToM is mainly tested on ToMi and a specific subset of BigToM questions that resemble those in ToMi. Similarly, the graph construction pipeline in SymbolicToM is designed specifically for ToMi. In contrast, our method, PercepToM, demonstrates flexibility and effectiveness when applied to data of varying formats, leading to significantly better generalizability compared to the existing methods. Although we do not include these two baselines in our main experiments, we present performance comparisons between PercepToM and these methods in Appendices F and G.
@@ 6 Conclusion
Inspired by the psychology literature, we evaluated the precursory inferences for human theory of mind (ToM) in large language models (LLM), aiming to broaden our insight into their ToM capabilities. To this end, we constructed perception-augmented ToM benchmarks, Percept-ToMi and Percept-FANToM, by annotating character perceptions about the contexts. Through evaluations and analyses of eight state-of-the-art LLMs, we found that they perform reasonably well in inferring others’ perceptions but struggle with inferring others’ beliefs based on that perceptual information. Based on these findings, we proposed a new framework, PercepToM, to improve LLM’s ToM reasoning. Our framework leverages LLMs’ strength in perception inference and enhances their perception-to-belief inference by extracting the relevant contexts. We expect our work to provide insights and enable further in-depth studies into the extent of LLMs’ ToM capabilities and targeted improvements in their weaknesses.
@@ 7 Limitations
We compare LLMs’ ToM performances between true belief and false belief scenarios, but not those between the different orders of ToM questions (e.g., first-order and second-order). Since higher-order ToM requires more inference steps, it will also be interesting to examine the differences in model behavior and capability in solving different orders of ToM questions in future work.
@@ 8 Societal and Ethical Considerations
Our use of FANToM dataset is consistent with its intended use, which is evaluation. We have adhered to the licenses of the benchmarks, ToMi and FANToM, in processing them to create our benchmarks, Percept-ToMi and Percept-FANToM. We plan to make our benchmarks publicly available with the license of Attribution-Noncommercial 4.0 International (CC BY-NC 4.0), allowing the sharing and adapting of the material.
@@
Acknowledgment
We thank Amazon for their gift in support of research on theory of mind. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00443251, Accurate and Safe Multimodal, Multilingual Personalized AI Tutors).
@@ A.1 Manual Verification of Perception Annotation in Percept-ToMi
SymbolicToM identifies the perceivers of a scene in ToMi deterministically based on a graphical representation of the world state constructed from a templated description of the scene. However, since it determines the perceiver of a scene as all entities in the same connected component based on the world graph, it produces two types of errors in the perceiver information, as shown in Table 3. By reviewing 50 samples of SymbolicToM output, we discover these error types and correct every occurrence of them throughout our dataset.
The perceiver of distractor sentences in ToMi, which describe a character’s opinion about an object, should be the character themselves, since they do not express their opinion to others in the scenario. However, the SymbolicToM-generated perceiver annotation includes other characters located in the same space. We therefore correct the perceivers for all such distractor sentences. Another error in perceiver annotations occur in the sentences preceding the location-disambiguating sentence, which specifies object locations, where their perceivers are annotated with ‘none.’ We align the perceiver annotations of these sentences with those of the subsequent location-disambiguating sentenc, since they are always paired and have the same perceivers.
@@
A.2 Perception Annotation Criteria for Percept-FANToM
The following criteria are used to determine the joining and leaving times of a character in a conversation within the FANToM dataset.
@@ A.2 Perception Annotation Criteria for Percept-FANToM
Appendix B Dataset Statistics
Table 4 presents a comparison of data statistics between perception-augmented ToM benchmarks and their corresponding source benchmarks.
Datasets True Belief False Belief
# Ctx. # Q. Avg. # Q. per Ctx. # Ctx. # Q. Avg. # Q. per Ctx.
P.I. PtoB ToM P.I. PtoB ToM
Disambiguated ToMi - Test Set
2793 2793 - - 1.0 1210 1210 - - 1.0
Percept-ToMi 300 1802 4.0 1.0 1.0 300 1804 4.0 1.0 1.0
FANToM 402 3432 - - 8.5 642 8530 - - 13.3
Percept-FANToM 340 13161 24.2 7.3 7.3 539 25279 23.3 11.8 11.8
Table 4: Comparison of dataset statistics between the source datasets and our proposed datasets. Number of contexts, the total number of questions, and the average number of questions per context for each task are compared. Our benchmarks expand upon the sampled contexts from the source datasets by incorporating two precursory inference tasks for Theory of Mind (ToM)—perception inference and perception-to-belief inference.
Appendix C Prompt Examples
@@ C.1 Perception Inference
C.2 Perception-to-Belief Inference
@@
Appendix D Input and Output Examples of PercepToM Pipeline
@@
[{"Ella likes the suit.": ["Ella"]},
@@
Here are the past scenes in sequence that Lucas knows about.
@@
Here are the past utterances in sequence that Gianna is aware of.
@@
Appendix E LLM Performances on Percept-ToMi and Percept-FANToM
@@
Appendix F Performance Comparison Between PercepToM and SymbolicToM
@@
Appendix G Performance Comparison Between PercepToM and SimToM
Model Method True Belief False Belief
GPT-4 Turbo SimToM 0.657 0.873
PercepToM 0.824 1.000
GPT-4o SimToM 0.797 0.450
PercepToM 0.659 0.915
Llama-3 70B Inst. SimToM 0.644 0.770
PercepToM 0.713 0.744
Mixtral 8x22B Inst. SimToM 0.677 0.660
PercepToM 0.727 0.964
Table 7: Performance comparison between SimToM and PercepToM on Fixed and Disambiguated ToMi (Sclar et al., 2023). Overall, PercepToM shows more robust performance across different models in both of true and false belief scenarios.
Table 7 shows the performances of PercepToM and SimToM on ToMi.